i(2+7i)(2-7i) write in standard form

Answer by AnlytcPhil(1806)

We need to know that iČ = -1

i(2 + 7i)(2 - 7i) =

(2i + 7iČ)(2 - 7i) =

[2i + 7(-1)](2 - 7i) =

(2i - 7)(2 - 7i) =

4i - 14iČ - 14 + 49i =

53i - 14iČ - 14 =

53i - 14(-1) - 14 =

53i + 14 - 14 =

53i

Edwin